OpenStudy (anonymous):

can anyone do this? 3x+1/14 = 2x/8

OpenStudy (anonymous):

what is the least common denominator of 14 and 8?

OpenStudy (anonymous):

2?

OpenStudy (mazmaz):

no

OpenStudy (anonymous):

Why?

OpenStudy (mazmaz):

because 2 cannot be divided by either of them you have found a common factor

OpenStudy (anonymous):

If you mean why as in why find the lcm then so you can get rid of the fractions

OpenStudy (anonymous):

112 then?

OpenStudy (mazmaz):

It is %^ because 56 div. by 14 is 4 and 56 div. by 8 is 7

OpenStudy (anonymous):

112 was good enough. You really just needed at common denominator. Now multiply both sides of the equation by 56. What do you get?

OpenStudy (anonymous):

14 * 56 = 784

OpenStudy (anonymous):

and 56 * 8 = 448

OpenStudy (anonymous):

no I mean like this 56(3x+1/14)=56(2x/8)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

So what is it?

OpenStudy (anonymous):

i have no idea

OpenStudy (mathmale):

@Thelazyone0305 : are you comfortable with our decision that the LCD is 56? Note that both 14 and 8 divide into 56 with no remainder in either case. Multiply all terms of 3x+1/14 = 2x/8 by 56. Cancel wherever possible:\[56(3x)+ 56(\frac{ 1 }{ 14 })=56\frac{ 2x }{ 8 }\] Doing this will eliminate the fractions. Your result? Type it in here, please.

OpenStudy (mathmale):

Sorry. That's 56(3x)+56(1/14)=56(2X/8)

OpenStudy (mathmale):

Your input, please?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

im only 12

OpenStudy (anonymous):

What is 56*3, 56*1/14, and 56*2/8? BTW I'm only 11 myself.

OpenStudy (mathmale):

Yes, and I'm only 71 myself. ;) Come on. You can multiply 56 by 3. Do this, please.

OpenStudy (mathmale):

Next, divide 56 by 14. Next, divide 56 by 8.